A mouse ran outside its home to get food. The cheese was located 3 meters north and it takes 20 seconds to reach it. The mouse went back running faster than before and went back home 3 meters south which took 10 seconds. What is the velocity of the mouse?
Answer by Alan3354(69443)

A mouse ran outside its home to get food. The cheese was located 3 meters north and it takes 20 seconds to reach it. The mouse went back running faster than before and went back home 3 meters south which took 10 seconds. What is the velocity of the mouse?
----------------
If you mean the average speed:
He ran 6 meters in 30 seconds --> 0.2 m/sec
================================
Going speed is 3/20 = 0.15 m/sec
returning is 3/10 = 0.3 m/sec
---
Avg = 2*0.15*0.3/(0.15+0.3) = 0.09/0.45 = 0.2 m/sec